# Quasi Zigzag Persistence: A Topological Framework for Analyzing Time-Varying Data

### v2

@@ 1 Introduction
Time varying data analysis Hamilton (1994); Box & Jenkins (1976) has been a fundamental challenge in the machine learning community for a long time, ranging from traditional univariate and multivariate time-series data to more complex structure such as sequence of graphs or point clouds. While the traditional time-series data has been handled effectively Wu et al. (2020), increasing complexity of modern applications has necessitated novel methodologies. Spatiotemporal Graph Neural Networks Yu et al. (2018); Oreshkin et al. (2021); Kan et al. (2022); Chu et al. (2023) have emerged as powerful tools to analyze graph sequences by decomposing the problem into two components: spatial dependency and temporal evolution. Similarly, specialized architectures have been proposed to process sequences of point clouds Liu et al. (2019); Fan et al. (2021); Huang et al. (2021); Rempe et al. (2020). However, most of these approaches utilize the geometric information, which is mostly local, among other information potentially missing crucial global patterns in the data. This becomes apparent in scenarios where the overall shape and connectivity of the data carry significant meaning, such as in brain connectivity patterns. Topological methods offer a compelling solution by capturing these global structures that persist across different scales. By augmenting the models with topological information, we can develop models that leverage both local geometric patterns and global topological characteristics for improved performance.
@@
Recently, topological Data Analysis (TDA) has emerged as a prominent field that can leverage topological information hidden behind the data into its analysis. Persistent Homology (PH), a cornerstone of TDA, provides a succinct method to extract and represent multiscale topological information. This capability has proved transformative in enhancing machine learning models, particularly where topological properties carry relevant information  Corbet et al. (2019); Chen et al. (2019); Carrière et al. (2020); Kim et al. (2020); Gabrielsson et al. (2020); Hofer et al. (2020); Zhao et al. (2020); Swenson et al. (2020); Carrière & Blumberg (2020); Vipond (2020); Bouritsas et al. (2022); Horn et al. (2022); Cang & Wei (2017); Zhang et al. (2022); Bauer et al. (2023); Liu et al. (2022); Lesnick & Wright (2015). Time-varying data needs extensions of the classical one-parameter persistence homology along two directions. First, the temporal component of the data necessitates another parameter, thus bringing multi parameter persistence homology (MPH) Botnan & Lesnick (2023) into picture. Second, monotone filtrations usually associated with the standard PH computations cannot accommodate deletions that are required for processing time-varying data. This necessitates to use zigzag persistent homology (ZPH) Carlsson & de Silva (2010) which has been shown to capture dynamic topological features in time-series data Myers et al. (2023); Chen et al. (2021); Tinarrage et al. (2025); Hacquard & Lebovici (2024); Coskunuzer et al. (2024); Beltramo et al. (2022). In effect, we need a combination of MPH and ZPH which allows PH (standard filtration) along one parameter and ZPH along the other. This requires changes to the structure of the underlying partially-ordered set (poset) to what we call quasi-zigzag poset resulting into Quasi Zigzag Persistent Homology (QZPH) introducing additional complexity and a new set of challenges.
@@
While MPH captures richer topological information than PH, it presents a different set of challenges due to lack of a complete invariant. This has motivated TDA researchers to look for different invariants that, although not complete, capture sufficient topological information  Vipond (2020)Corbet et al. (2019); Carrière & Blumberg (2020); Russold & Kerber (2024); Scoccola et al. (2024); Loiseaux et al. (2023); Xin et al. (2023); Mukherjee et al. (2024). The juxtaposition of ZPH to MPH needs the zigzag generalizations of these methods.
@@
In this paper, we find that, one of these methods called Gril Xin et al. (2023) that computes a landscape function Bubenik (2015) using generalized ranks of intervals Kim & Mémoli (2021) adapts naturally to the framework of QZPH both conceptually and computationally. This adaptation leads to our key contribution which we call Zz-Gril, a new topological invariant that extends Gril Xin et al. (2023) framework to capture multiscale topological information in time-evolving data. On the theoretical front, we prove the stability of Zz-Gril and show that the generalized rank of a specific type of subposet, in the setting of QZPH, can be computed by an algorithm presented in Dey et al. (2024). We use this to devise an algorithm to compute Zz-Gril, making this method practical for real-world applications. We demonstrate the practical value of Zz-Gril by experimenting on data from critical healthcare monitoring tasks, such as sleep-stage detection via ECG and multivariate time-series data from other applications. We augment the topological information captured by Zz-Gril with existing machine learning models and show that the addition of topological information improves the performance of the models as exhibited by experimental results in Section 5  [^1] Code is available at https://github.com/TDA-Jyamiti/zzgril.
@@
In a very recent paper Flammer & Hüper (2024), the authors consider a similar setup and develop a method to visualize such data using spatiotemporal persistence landscapes. We acknowledge that the setup is similar but there are many differences in the overall results. One main difference is that Flammer & Hüper (2024) uses rectangles as intervals that allow them to use a simple modification of a result in Dey et al. (2024) whereas we use more general subposets for which we prove Theorem 4.5, which allows us to compute generalized ranks with zigzag modules. Because of the same reason, the overall algorithm to compute the landscape also becomes different. Additionally, we give an efficient algorithm to build a quasi zigzag bi-filtration from an input time-varying data. Finally, we develop a complete ML pipeline that we experiment with real world time-varying data which is not considered in Flammer & Hüper (2024).
@@ 2 Overview
We begin by recalling the definition of a simplicial complex. Given a finite vertex set $V$, a simplicial complex $K=K(V)$ is a collection of subsets of $V$ such that if a subset $\sigma\subseteq V$ is in $K$, then all subsets $\tau\subset\sigma$ are also in $K$. Each element in $K$ with cardinality $k+1$ is called a $k$-simplex or simply a simplex. A (non-zigzag) filtration $\mathcal{F}$ is a nested sequence of simplicial complexes indexed by integers/natural numbers: $\mathcal{F}:K_{0}\hookrightarrow K_{1}\hookrightarrow\cdots\hookrightarrow K_{n}$ where the inclusion arrows are all in forward direction. If these arrows are not all in forward direction, we get a filtration called zigzag filtration
$\mathcal{Z}:K_{0}\hookrightarrow K_{1}\hookleftarrow\cdots\hookrightarrow K_{n}.$
@@
From such a quasi zigzag bi-filtration, we obtain a so-called persistence module by considering the simplicial homology group $H_{p}(K_{i})$ in some dimension $p$ for every simplicial complex $K_{i}$. The inclusion maps in zigzag or non-zigzag filtrations translate to homomorphisms among these homology groups which become linear maps when homology groups turn into vector spaces under a field coefficient, say $k$. These vector spaces with dimension, say $d$, are often depicted with their isomorphic counterparts of $d$ copies of $k$, written as $k^{d}$. We refer to such a persistence module as a quasi zigzag persistence module; In Figure 1 (right), we show the quasi zigzag persistence module for zeroth homology group $H_{0}(\cdot)$ that represents the number of components.
The indexing set of a non-zigzag, zigzag, or quasi zigzag filtration can be thought of as a poset $P$ where for every pair of comparables $\mathbf{p}\leq\mathbf{q}$ in $P$, one has the inclusion of complexes $K_{\mathbf{p}}\subseteq K_{\mathbf{q}}$. A non-zigzag poset can be indexed by integers in $\mathbb{Z}$. A zigzag poset that arises in our case is a subposet of the poset $\mathbb{ZZ}$ defined below.
Definition 2.1 ((Quasi) Zigzag Poset).
Let $\mathbb{ZZ}$ be the subposet of $\mathbb{Z}^{\mathsf{op}}\times\mathbb{Z}$ defined as
$\mathbb{ZZ}\coloneqq\{(i,j):i\in\mathbb{Z}^{\mathsf{op}},j\in\{i,i-1\}\}.$
A poset $P$ is called zigzag if $P$ is a subposet of $\mathbb{ZZ}$ and is quasi zigzag if it is a subposet of the product $\mathbb{ZZ}\times\mathbb{Z}$ with the product order.
Now, we formally define persistence modules over different kind of posets.
@@
A persistence module $M$ over a poset $P$ is a functor $M:P\rightarrow{\bf vec}$ where $P$ is considered as a category and ${\bf vec}$ is the category of finite dimensional vector spaces. Stating more explicitly, $M$ is a collection of vector spaces $\{M_{\mathbf{p}}\}_{\mathbf{p}\in P}$ and linear maps $\{M_{\mathbf{p}\rightarrow\mathbf{q}}\}$ for every comparable $\mathbf{p}\leq\mathbf{q}$ in $P$. It is called a zigzag module if $P$ is zigzag and a quasi zigzag module if $P$ is quasi zigzag.
Next, we introduce the concepts of connectedness and intervals in posets which will be used for further exposition. We say a poset $P$ is connected if there is a sequence of points $p=r_{1},\ldots,r_{t}=q$ in $P$ so that $r_{i}\leq r_{i+1}$ or $r_{i+1}\leq r_{i}$ for every $i\in\{1,\ldots,t-1\}$. We say a subposet $I\subseteq P$ is convex if for every pair $p,q\in I$ and every $p\leq r\leq q$, we have $r$ in $I$. An interval $I$ in a poset $P$ is a subposet which is both convex and connected. An interval $I$ with $m\geq 0$ minima and $n\geq 0$ maxima is called an $(m,n)$-interval. In the special case of $\mathbb{Z}^{2}$, every interval has a special form where the sets of minima and maxima along with other points constitute staircases that can be visualized as part of the boundary of the interval; see Figure 2.
The main construct upon which we build Zz-Gril involves the so-called generalized ranks  Kim & Mémoli (2021) which can be computed efficiently by a zigzag persistence algorithm Dey et al. (2024). Let $P$ be a quasi zigzag poset over which a quasi zigzag persistence module $M$ is defined and $I\subseteq P$ be a subposet. When $M$ is induced by a bi-filtration in QZPH, the generalized rank of the module restricted to $I$ measures the multiplicity of the homological features (# of independent homological classes) that have support over the entire poset $I$. For example, considering $I=P$ (shaded light green), we see that, in Figure 1(middle), one component survives the entire poset. Then, its generalized rank for zero-th homology $H_{0}$ is 1. Similarly, the subposet in the bottom row (light blue) has a generalized rank of 3 for $H_{0}$ and the subposet in bottom right square (light yellow) has a generalized rank of 2 for $H_{0}$. We formally define generalized rank below.
@@
Let $M:P\to\mathbf{vec}$ be a persistence module from a finite and connected poset $P$. The restriction of $M$ to a subposet $I$ of $P$, $M|_{I}$, is the collection of vector spaces $M_{\mathbf{p}}$ for $\mathbf{p}\in I$ along with linear maps $M_{\mathbf{p}\to\mathbf{q}}$ for all comparable $\mathbf{p}\leq\mathbf{q}$ in $I$. Then, the generalized rank of $M$ over $I$ is defined as the rank of the canonical linear map from $\mathsf{lim}M|_{I}$ to $\mathsf{colim}M|_{I}$
@@
One of the very useful properties of generalized rank is that it is monotone w.r.t. the inclusion of posets. We use this property to define Zz-Gril.
Monotone Property of Generalized Rank. It follows from the definition that $\mathsf{rk}^{M}(I)\leq\mathsf{rk}^{M}(J)$ for all $J\subseteq I$, where $I\text{ and }J$ are subposets in $P$.
To exploit the topological features captured by generalized ranks, we use a cover of the quasi zigzag poset with a specific set of subposets called worms. When endowed with $\mathbb{Z}^{2}$ ordering, these subposets become $(3,3)$-intervals around a set of chosen centers with some chosen widths. Definition 3.1 gives precise definitions of these terms. We expand each of these worms, that is, increase the widths of these $(3,3)$-intervals while keeping the centers fixed. The monotonicity of the generalized rank ensures that it can only decrease or remain the same by this expansion. In Figure 1(right) a $(1,1)$-interval (rectangle) is expanded maximally to decrease the generalized rank from $3$ to $1$. The width for which the rank drops below a chosen threshold is taken as the Zz-Gril value at the center of the interval. Defintion 3.2 makes this concept precise. The set of Zz-Gril values at the chosen centers makes a vector that we use in the learning pipeline.
@@ 3 Zz-Gril
In this section, we introduce concepts and definitions required for Zz-Gril. Then, we define Zz-Gril and discuss its theoretical properties. Keep in mind that, even if we do not explicitly state it always, these concepts apply to QZPH setting where vector spaces and linear maps in question are given by simplicial homology groups and homomorphisms between them induced by the input bi-filtrations.
Consider the poset $\mathbb{ZZ}\times\mathbb{Z}$ with the product order, i.e., $(\mathbf{z_{1}},z_{2})\leq(\mathbf{w_{1}},w_{2})$ if $\mathbf{z_{1}}\leq\mathbf{w_{1}}$ in $\mathbb{ZZ}$ and $z_{2}\leq w_{2}$ in $\mathbb{Z}$. Note that $\mathbb{ZZ}\times\mathbb{Z}$ is equivalent to $\mathbb{Z}^{2}$ as sets. Thus, every subposet $P\subseteq\mathbb{ZZ}\times\mathbb{Z}$ can be thought as a subposet $P^{\mathbb{Z}^{2}}\subseteq\mathbb{Z}^{2}$ that is endowed with $\mathbb{Z}^{2}$ ordering. Interestingly, an interval in $\mathbb{Z}^{2}$ may not remain an interval in $\mathbb{ZZ}\times\mathbb{Z}$ and vice-versa.
We define a quasi zigzag bi-filtration over a quasi zigzag poset $P$ as a collection of simplicial complexes $\{K_{\mathbf{p}}\}_{\mathbf{p}\in P}$, where $K_{\mathbf{p}}\subseteq K_{\mathbf{q}}$ for all comparable $\mathbf{p}\leq\mathbf{q}$. Akin to 1-parameter and zigzag persistence modules, inclusion-induced linear maps between the homology vector spaces of simplicial complexes in a quasi zigzag bi-filtration define a quasi zigzag persistence module.
@@
Let $\mathbf{p}\in\mathbb{ZZ}\times\mathbb{Z}$ and $\delta\in\mathbb{Z}$ be given. Let $\boxed{\mathbf{p}}_{\delta}$ denote the $\delta$-square centered at $\mathbf{p}$, i.e., $\boxed{\mathbf{p}}_{\delta}\coloneqq\{\mathbf{z}\in\mathbb{ZZ}\times\mathbb{Z}:||\mathbf{p}-\mathbf{z}||_{\infty}\leq\delta\}$. Then, a worm centered at $\mathbf{p}$ with width $\delta$ is defined as the union of $\boxed{\mathbf{p}}_{\delta}$ with the two $\delta$-squares $\boxed{\mathbf{q}}_{\delta}$ centered at points $\mathbf{q}=\mathbf{p}\pm(\delta,-\delta)$ on the off-diagonal line segment. We denote the worm as $\boxed{\mathbf{p}}_{\delta}^{2}$. The superscript denotes the number of $\delta$-squares in the union apart from $\boxed{\mathbf{p}}_{\delta}$.
We note that computing $||\cdot||_{\infty}$ in $\mathbb{ZZ}\times\mathbb{Z}$ is equivalent to computing it in $\mathbb{Z}^{2}$ using the notion of set equivalence. We can see that a worm turns out to be a $(3,3)$-interval in $\mathbb{Z}^{2}$. Refer to Figure 2 for an illustration of a worm.
@@
The basic idea of Zz-Gril is similar to Gril. However, we note that the underlying poset structure is very different. We cover $\mathbb{ZZ}\times\mathbb{Z}$ with subposets of specific shapes (worms), and compute generalized rank over these worms to define the landscape function (Zz-Gril). Next, we show that Zz-Gril as an invariant remains stable.

### v3

@@ 1 Introduction
Time varying data analysis Hamilton (1994); Box & Jenkins (1976) has been a fundamental challenge in the machine learning community, ranging from traditional time-series data to more complex structures such as sequences of graphs or point clouds. While traditional time-series have been handled effectively Wu et al. (2020), the complexity of modern applications necessitates novel methodologies. Spatiotemporal Graph Neural Networks Yu et al. (2018); Oreshkin et al. (2021); Kan et al. (2022); Chu et al. (2023) and specialized architectures for point cloud sequences Liu et al. (2019); Fan et al. (2021); Huang et al. (2021); Rempe et al. (2020) have emerged as powerful tools. However, most of these approaches utilize local geometric information, potentially missing crucial global patterns. This is particularly apparent where the overall shape carries significant meaning, such as in brain connectivity patterns. Topological methods offer a compelling solution by capturing these global, multi-scale structures. By augmenting models with topological information, we can leverage both local geometric patterns and global characteristics for improved performance.
@@
Recently, Topological Data Analysis (TDA) has become a prominent field for leveraging such hidden information. Persistent Homology (PH), a cornerstone of TDA, provides a succinct method to extract multiscale topological features. This has been transformative in enhancing machine learning models. Analyzing time-varying data requires extending classical PH in two directions. First, the temporal component necessitates an additional parameter, bringing multiparameter persistence homology (MPH) Botnan & Lesnick (2023) into the picture. Second, standard PH computations use monotone filtrations which cannot accommodate the deletions required for time-varying data. This calls for zigzag persistent homology (ZPH) Carlsson & de Silva (2010), which captures dynamic topological features in time-series data Myers et al. (2023); Chen et al. (2021); Tinarrage et al. (2025); Hacquard & Lebovici (2024); Coskunuzer et al. (2024); Beltramo et al. (2022). Effectively, we need a combination of MPH and ZPH, allowing standard PH filtration along one parameter and ZPH along another. This requires structuring the underlying partially-ordered set (poset) as a quasi-zigzag poset, leading to what we term Quasi Zigzag Persistent Homology (QZPH), which introduces a new set of challenges.
@@
While MPH captures rich information, it suffers from the lack of a complete invariant, motivating the search for other informative, though incomplete, invariants Vipond (2020); Corbet et al. (2019); Carrière & Blumberg (2020); Scoccola et al. (2024); Loiseaux et al. (2023); Russold & Kerber (2024); Xin et al. (2023); Mukherjee et al. (2024). The application of ZPH to MPH requires zigzag generalizations of these methods.
@@
In this paper, we find that one such method, Gril Xin et al. (2023) , which computes a landscape function Bubenik (2015) using generalized ranks of intervals Kim & Mémoli (2021), adapts naturally to the QZPH framework. This adaptation leads to our key contribution, Zz-Gril, a new topological invariant that extends the Gril framework to capture multiscale topological information in time-evolving data. On the theoretical front, we prove the stability of Zz-Gril and show that the generalized rank over a specific type of subposet can be computed efficiently using an algorithm from Dey et al. (2024). This allows us to devise a practical algorithm for computing Zz-Gril. We demonstrate its value by augmenting machine learning models for tasks like sleep-stage detection from ECG and action classification from multivariate time-series, showing improved performance (Section 5).  [^1] Code is available at https://github.com/TDA-Jyamiti/zzgril.
@@
A recent paper Flammer & Hüper (2024) considers a similar setup for visualizing data with spatiotemporal persistence landscapes. While the setup is similar, our work differs significantly. Notably, Flammer & Hüper (2024) use simple rectangular intervals, allowing a direct modification of a result in Dey et al. (2024). In contrast, we use more general subposets, for which we prove Theorem 4.6, enabling the computation of generalized ranks for zigzag modules and leading to a different overall algorithm. We also provide an efficient algorithm to construct a quasi-zigzag bi-filtration from raw time-varying data and develop a complete ML pipeline that we validate on real-world datasets not considered in Flammer & Hüper (2024).
@@ 2 Overview
We begin with essential topological concepts. A simplicial complex $K$ is a collection of subsets (simplices) of a finite vertex set $V$, such that if a simplex $\sigma$ is in $K$, then all its subsets are also in $K$. A (non-zigzag) filtration $\mathcal{F}$ is a nested sequence of simplicial complexes indexed by integers: $\mathcal{F}:K_{0}\hookrightarrow K_{1}\hookrightarrow\cdots\hookrightarrow K_{n}$. If some inclusions are reversed, we get a zigzag filtration, e.g., $\mathcal{Z}:K_{0}\hookrightarrow K_{1}\hookleftarrow\cdots\hookrightarrow K_{n}$.
@@
The indexing set of a filtration is a partially ordered set (poset). A non-zigzag filtration can be indexed by the integers $\mathbb{Z}$, while the quasi zigzag bi-filtration is indexed by a product poset $\mathbb{ZZ}\times\mathbb{Z}$, where $\mathbb{ZZ}$ denotes a zigzag poset that arises in our case(see Definition 3.1). Applying simplicial homology functor (e.g., $H_{p}$) to each complex in the filtration yields a persistence module: a collection of vector spaces connected by linear maps according to the inclusions. For a quasi zigzag bi-filtration, this results in a quasi zigzg persistence module. In Figure 1 (right), we show the quasi zigzag persistence module for zeroth homology group $H_{0}(\cdot)$ that represents the number of components.
@@
A persistence module over a poset $P$ is a functor $M\colon P\to\mathbf{vec}$, i.e., a collection of vector spaces $\{M_{\mathbf{p}}\}_{\mathbf{p}\in P}$ along with linear maps $M_{\mathbf{p}\to\mathbf{q}}\colon M_{\mathbf{p}}\to M_{\mathbf{q}}$ for every $\mathbf{p}\leq\mathbf{q}$ in $P$. Here, $\mathbf{vec}$ denotes the category of finite dimensional vector spaces.
Next, we introduce the concept of intervals. A finite poset $P$ is connected if for every pair $p,q\in P$ there is a sequence of points $p=r_{1},\ldots,r_{t}=q$ in $P$ so that $r_{i}\leq r_{i+1}$ or $r_{i+1}\leq r_{i}$ for all $i\in\{1,\ldots,t-1\}$. An interval $I\subseteq P$ is a connected subposet that is convex w.r.t. poset order, i.e., if $p,q\in I$ and $p\leq r\leq q$, then $r\in I$. Figure 2 depicts an interval in $\mathbb{Z}^{2}$.
Our framework relies on the generalized rank  Kim & Mémoli (2021) of a persistence module over a subposet $I\subseteq P$, which can be computed for $2$-parameter persistence modules with the algorithm Dey et al. (2024). The generalized rank of the module restricted to $I$ measures the multiplicity of the homological features (# of independent homological classes) that have support over the entire poset $I$. For example, in Figure 1, the generalized rank, for $H_{0}$, over the entire poset (light green) is 1 as one component survives throughout, while the generalized rank over the bottom row (light blue) is 3 and that over the bottom right square (light yellow) is 2.
@@
Let $M:P\to\mathbf{vec}$ be a persistence module, where $P$ is a finite connected poset. Let $M|_{I}$ denote the restriction of $M$ to a subposet $I$ of $P$. Then, the generalized rank of $M$ over $I$ is defined as the rank of the canonical linear map from $\mathsf{lim}M|_{I}$ to $\mathsf{colim}M|_{I}$
@@
A key property of generalized rank is its monotonicity: $\mathsf{rk}^{M}(I)\leq\mathsf{rk}^{M}(J)$ for all $J\subseteq I$, where $I\text{ and }J$ are subposets in $P$. We use this to define Zz-Gril.
To extract the topological information from a quasi zigzag persistence module, we cover the quasi zigzag poset with a specific type of subposets called worms. A worm has a notion of a center and a width. Definition 3.3 gives precise definitions of these terms. We expand each of these worms, i.e., increase the widths while keeping the centers fixed. The monotonicity of the generalized rank ensures that the rank can only decrease by this expansion. In Figure 1(right) a subposet (rectangle) is expanded maximally to decrease the generalized rank from $3$ to $1$. The width for which the rank drops below a chosen threshold is taken as the Zz-Gril value. Defintion 3.4 makes this concept precise. The set of Zz-Gril values at the chosen centers makes a vector that captures the topological information in the quasi zigzag persistence module.
@@ 3 Zz-Gril
In this section, we introduce concepts and definitions required for Zz-Gril. Then, we define Zz-Gril and discuss its theoretical properties.
Definition 3.1 (Zigzag poset).
Zigzag poset $\mathbb{ZZ}$ is defined as a subposet of $\mathbb{Z}^{\mathsf{op}}\times\mathbb{Z}$ given by
$\mathbb{ZZ}\coloneqq\{(i,j)\colon i\in\mathbb{Z},j\in\{i,i-1\}\},$
where $\mathbb{Z}^{\mathsf{op}}$ denotes the opposite poset of $\mathbb{Z}$.
Consider the poset $\mathbb{ZZ}\times\mathbb{Z}$ with the product order, i.e., $(\mathbf{z_{1}},z_{2})\leq(\mathbf{w_{1}},w_{2})$ if $\mathbf{z_{1}}\leq\mathbf{w_{1}}$ in $\mathbb{ZZ}$ and $z_{2}\leq w_{2}$ in $\mathbb{Z}$. Note that $\mathbb{ZZ}\times\mathbb{Z}$ is equivalent to $\mathbb{Z}^{2}$ as sets. Thus, every subposet $I\subseteq\mathbb{ZZ}\times\mathbb{Z}$ can be thought as a subposet $I^{\mathbb{Z}^{2}}\subseteq\mathbb{Z}^{2}$ that is endowed with $\mathbb{Z}^{2}$ ordering. Observe that an interval in $\mathbb{Z}^{2}$ may not remain an interval in $\mathbb{ZZ}\times\mathbb{Z}$ and vice-versa.
We define a quasi zigzag bi-filtration as a collection of simplicial complexes $\{K_{\mathbf{p}}\}_{\mathbf{p}\in\mathbb{ZZ}\times\mathbb{Z}}$, where $K_{\mathbf{p}}\subseteq K_{\mathbf{q}}$ for all comparable $\mathbf{p}\leq\mathbf{q}$. Let $\mathbf{vec}$ denote the category of finite dimensional vector spaces.
Definition 3.2 (Quasi zigzag persistence module).
A persistence module $M\colon\mathbb{ZZ}\times\mathbb{Z}\to\mathbf{vec}$ is called a quasi zigzag persistence module.
@@
[…]
We note that computing $||\cdot||_{\infty}$ in $\mathbb{ZZ}\times\mathbb{Z}$ is equivalent to computing it in $\mathbb{Z}^{2}$ using the notion of set equivalence. We can see that a worm turns out to be an interval in $\mathbb{Z}^{2}$. Refer to Figure 2 for an illustration of a worm.
@@
The basic idea of Zz-Gril is similar to Gril. However, we note that the underlying poset structure is very different. For computations, we consider a finite subposet of $\mathbb{ZZ}\times\mathbb{Z}$ and cover it with worms. Then, we compute generalized rank over these worms to define the landscape function (Zz-Gril).